OpenStudy (anonymous):

Madeline and Jonathan want to purchase a home in seven years. They will contribute $720 each month to a savings account with 6.35% interest, compounded twice a year. What is the future value of this investment, when Madeline and Jonathan need to make a down payment?

OpenStudy (anonymous):

$77,065.89 $12,449.06 $6,108.44 $74,694.35

OpenStudy (anonymous):

What math is this?

OpenStudy (anonymous):

Algebra with Financial Apps I know i'm supposed to use the future value formula, but i don't know which numbers to plug in where...

OpenStudy (anonymous):

contributions per two years = 720*24 = $17,280. After two years, due to interest the money in account = $17,280*(1.0635) = $18,337.28 After four years it is ($18,337.28 + $17,280)*(1.065) = $37,932.4032 After six years it is ($37,932.4032 + $17,280)*(1.065) = $58,801.21 After the seventh year it is $58,801.21+$17,280 = $77,081.21 and this should be the down payment amount.

OpenStudy (anonymous):

Well i am in Consumer math and i did take algebra but a different one sorry if i couldn't figure it out, i wish i really knew it.

OpenStudy (anonymous):

that's ok. And Thanks @abhilashkusam!

OpenStudy (anonymous):

:)